# Normal Quaternionic Matrices and Finitely Generated Witt Rings

### v1

@@ 3. The Abstract 2-Brauer Group
Let $S=(G,-1,q)$ be a quaternionic structure. Let $(B,\ast)$ be a group into which the set of quaternions $Q$ embeds such that
@@
Let $S=(G,-1,q)$ be a quaternionic structure with 2B-group $B$ and set of quaternions $Q$. The following hold:
@@
    $B$ is a group of exponent $2$ with neutral element 00.
@@
(a): By definition, we have $q(a,b)\ast q(a,b)=q(a,b^{2})=q(a,1)=0$ and $q(a,b)\ast 0=q(a,b)\ast q(a,1)=q(a,b)$, so $B$ is a group of exponent $2$ with neutral element 00.
@@
so $B$ is generated by the $q(a_{i},a_{j})$. By symmetry (Q2), we can always assume $i\leq j$ here. Furthermore, choosing $a_{0}=-1$ if $-1\not=1\in G$, for any $i\in\{0,\ldots,n-1\}$, we get (cf. Remark 1.1)
@@
Hence, $B$ is generated by $(q(a_{i},a_{j})\mid 0\leq i<j\leq n-1)$ and, if $1$-1$\not$=1- 1 ≠ 1 additionally $q(a_{0},a_{0})$. The assertion follows. ∎
@@
Let $S=(G,-1,q)$ and $S^{\prime}=(G^{\prime},-1^{\prime},q^{\prime})$ be quaternionic structures with full 2B-groups $B$ and $B^{\prime}$. The following hold:
@@
If $F$ is a field and $I$ the fundamental ideal of its Witt ring, the factor group $I^{2}/I^{3}$ is isomorphic to the $2$-torsion part of the Brauer group by Merkurjev’s Theorem [Merkurjev_KTheory]. In our abstract setting, the full $2B$-group will function as a substitute for $I^{2}/I^{3}$. We do so to avoid assuming $AP_{3}$ as an additional property of quaternionic structures. In the remainder of this section, we show that both $I^{2}/I^{3}$ and the full 2B-group are in fact the same for elementary types.
@@
For the reader’s convenience, we recall Marshall’s results from [Marshall_AbstractWittRings, proof of 5.7, 5.17] on how the Arason-Pfister property $AP_{k}$ behaves for direct products and group extensions. Note that the statement for the direct product is only proven there for $k\in\{1,2\}$ but the proof works in the same way for larger $k$.
@@ 4. Quaternionic Matrices
In the following, we only consider structures $S$ of finite order $n\in\mathbb{N}$. In this section, we establish a unique matrix $M_{S}\in\operatorname{Mat}_{n}(\mathbb{N}_{0})$ by which we can identify $S$. This will be the groundwork for our computations later on. We usually write $S=(G,-1,q)$ with set of quaternions $Q$ and full $2B$-group $B$. We call its dimension $\dim(S):=\dim(B)$, which is finite by Lemma 3.1(c), the dimension of $S$.
@@
For $n\in\mathbb{N}$, we often consider matrices $M=(M_{i,j})_{0\leq i,j\leq n-1}\in\operatorname{Mat}_{n}(\mathbb{N}_{0})$ where we start counting the indices for rows and columns at 00. We bestow the set $\operatorname{Mat}_{n}(\mathbb{N}_{0})$ with the lexicographical order $<_{\text{lex}}$, applied row by row, and extending the usual order on $\mathbb{N}_{0}$. We simply say that a matrix is ’smaller/greater/…’ than another if it is lexicographically smaller/greater/…
@@
defines a group of exponent $2$ $(\mathbb{N}_{0},\ast_{\mathbb{N}_{0}})$ where any set $\{0,1,\ldots,2^{n}-1\}$ defines a subgroup with basis $(2^{k}\mid 0\leq k\leq n-1)$. This is easily verified using linear algebra. We often simply write $\ast$ instead of $\ast_{\mathbb{N}_{0}}$ and define the completion of $M$ as the matrix
@@
We also define $(B(M),\overline{\ast})$ as the free group of exponent $2$ $\langle\overline{M}_{I,J}\mid I,J\subset\{0,\ldots,n-1\}\rangle$ with the (linkage) relations
@@
We say that a basis $\mathcal{A}=(a_{0},\ldots,a_{n-1})$ of $G$ is a basis of S if either $-1=1$ or $a_{0}=-1$. Given a basis $\mathcal{B}=(q_{0},\ldots,q_{m-1})$ of $B$, we define the group isomorphism
@@
If $\mathcal{B}$ is fixed, we often identify $(B,\ast)$ with ($(\{0,\ldots,2^{m}-1\},\ast)$, writing $q_{\mathcal{B}}(a,b)$ or – in abuse of notation – simply $q(a,b)$ instead of $\phi_{\mathcal{B}}(q(a,b))$.
@@
Note that the matrices defined as above are in general not the incidence matrices as studied by Kula (cf. [Kula_FinitelyGeneratedWittRings]). The explicit matrices given here are in fact the lexicographically minimal ones for the respective structures which we leave for the reader to check. The minimality will be of importance for defining a unique quaternionic matrix of $S$.
@@
    • if $S\cong\mathbb{L}_{2k,0}$: $M_{\mathcal{A},\mathcal{B}}$ is the matrix $E_{2k}\in M_{2k}(\mathbb{N}_{0})$ with entry $1$ on the anti-diagonal and entry 00 else.
@@
(b): The properties (b)(M2) and (b)(M4) follow directly from (Q2) and (Q4). For (b)(M3), it suffices to see that $q(a,a)=q(a,-1)$ for all $a\in G$ by Remark 1.1 and this equals 00 if $1=-1$ in $G$. Since we have chosen $a_{0}=-1$ in the other case, (b)(M3) follows.
@@
Let $M\in\operatorname{Mat}_{n}(\mathbb{N}_{0})$ be a prequaternionic matrix of type $\epsilon\in\{0,1\}$. Let $G$ be a group of exponent $2$ with neutral element $1$ and basis $\mathcal{A}:=(a_{0},\ldots,a_{n-1})$. If $\epsilon=0$, choose $-1:=1\in G$, else $-1:=a_{0}$. Define $Q$ as the set of entries of $\overline{M}$, naturally containing 0. Then
@@
defines a quaternionic structure $S(M)=(G,-1,q)$ with $s(S)=1$ if $\epsilon=0$ and $s(S)>1$ else. For any basis $\mathcal{B}$ of the full $2B$-group of $S(M)$ and $I,J\subset\{0,\ldots,n-1\}$, it satisfies
@@
(a): Write $\mathcal{A}$ and $\mathcal{A}^{\prime}$ for the bases of $S$ and $S^{\prime}$ as in Proposition 4.4. It is clear that the basis transformation $\mathcal{A}\mapsto\mathcal{A}^{\prime}$ defines an isomorphism of quaternionic structures as it preserves the $(-1)$-element ($M,M^{\prime}$ are of the same type) and the zero quaternions.
(b): As before, this follows from the fact that the canonical isomorphism of groups of exponent $2$ preserves $-1$ and the zero quaternions. ∎
@@
$(\mathbb{L}_{1}\times({\color[rgb]{0,1,0}\mathbb{L}_{1}}\times{\color[rgb]{0,0,1}\mathbb{L}_{1}}){\color[rgb]{1,0,0}\Delta}){\color[rgb]{1,.5,0}\Delta}=:S=:(G,-1,q)$
with square class group $G=\{1,a_{1}\}\times{\color[rgb]{0,1,0}\{1,b_{0}\}}\times{\color[rgb]{0,0,1}\{1,a_{2}\}}\times{\color[rgb]{1,0,0}\{1,a_{3}\}}\times{\color[rgb]{1,.5,0}\{1,a_{4}\}}$ according to construction and coloring. Then $a_{0}:=-1=a_{1}{\color[rgb]{0,1,0}b_{0}}{\color[rgb]{0,0,1}a_{2}}$. We compute $M:=M_{\mathcal{A}}(S)$ for the basis $\mathcal{A}=(a_{0},\ldots,a_{4})$. We have $q(a_{0},a)\not=0$ for all $a\in G\setminus\{1\}$ ($S$ is pythagorean), so all entries of the first row of $M$ must be independent, meaning that by (c)(M0), we have (apply Proposition 5.1 from the following section or a straightforward computation) the first row (1 2 4 8 16)124816(1\ 2\ 4\ 8\ 16)( 1 2 4 8 16 ). Furthermore, by the properties of the direct product $\mathbb{L}_{1}\times({\color[rgb]{0,1,0}\mathbb{L}_{1}}\times{\color[rgb]{0,0,1}\mathbb{L}_{1}}){\color[rgb]{1,0,0}\Delta}$, we get $q(a_{1},{\color[rgb]{0,0,1}a_{2}})=0=q(a_{1},{\color[rgb]{1,0,0}a_{3}})$. Applying (b)(M2) and (b)(M3) there are $4$ more values to compute:
@@
$\dim(S)=1+(({\color[rgb]{0,1,0}1}+{\color[rgb]{0,0,1}1})+{\color[rgb]{1,0,0}2})+{\color[rgb]{1,.5,0}4}=9.$
@@
Let $\epsilon\in\{0,1\}$. For any $n\in\mathbb{N}$, we have a bijection between the set of normal quaternionic matrices of type $\epsilon$ in $\operatorname{Mat}_{n}(\mathbb{N}_{0})$ and the isomorphy types of quaternionic structures $S$ of order $n$ with $-1=1$ (if $\epsilon=0$) or $1$-1$\not$=1- 1 ≠ 1 (if $\epsilon=1$), mapping $M$ to $S(M)$.
@@
Let $M\in\operatorname{Mat}_{n}(\mathbb{N}_{0})$ be a quaternionic matrix of type $\epsilon\in\{0,1\}$. We say that $M^{\prime}\in\operatorname{Mat}_{n}(\mathbb{N}_{0})$ is a simple transformation of $M$ if there are $i,j\in\{0,\ldots,n-1\}$, $i\not=j$ such that
@@
The simple transformations display basis transformations of quaternionic structures in the following way: Take a quaternionic structure $S=(G,-1,q)$ with basis $\mathcal{A}=(a_{0},\ldots,a_{n-1})$ and basis $\mathcal{B}$ of the full $2B$-group. As $G$ is an $\mathbb{F}_{2}$ vector space, we cannot scale $\mathcal{A}$ nontrivially, so if $\mathcal{A}^{\prime}$ is another basis of $S$, we can obtain it from $\mathcal{A}$ by an iteration of swapping elements $a_{i}\leftrightarrow a_{j}$ or replacing $a_{i}$ by $a_{i}a_{j}$. The following lemma shows that the matrix transformations defined above translate the effects of these basis transformations to the quaternionic matrices, going from $M_{\mathcal{A},\mathcal{B}}$ to $M_{\mathcal{A}^{\prime},\mathcal{B}}$ by an iteration of simple transformations. For the canonical choice of $\mathcal{B}$ (cf. Lemma 4.2(c)) in this setting, we have to consider reduced quaternionic matrices:
@@
(c): First, we point out that applying simple transformations $(i\leftrightarrow j)$ or $(i\mapsto i\ast j)$ just means swapping the $i$-th and $j$-th basis element in the induced basis of $S(M)$ or replacing $a_{i}$ by $a_{i}a_{j}$. Both are basis transformations of $S(M)$ as by the choice of $i,j$, the element $-1$ is fixed. In particular, simple transformations and by iteration also general transformations of $M$ are again quaternionic matrices of the same type that are isomorphic to $M$. By (a), reduction of quaternionic matrices also does not change the isomorphy type, so the ’if’ part of the statement holds.

### v2

@@ 3. The Abstract 2-Brauer Group
Let S=(G,1,q)S=(G,-1,q) be a quaternionic structure. Let $(B,\ast)$ be a group into which the set of quaternions $Q$ embeds such that
@@
Let S=(G,1,q)S=(G,-1,q) be a quaternionic structure with 2B-group $B$ and set of quaternions $Q$. The following hold:
@@
    $B$ is a group of exponent $2$ with neutral element 0.
@@
(a): By definition, we have $q(a,b)\ast q(a,b)=q(a,b^{2})=q(a,1)=0$ and $q(a,b)\ast 0=q(a,b)\ast q(a,1)=q(a,b)$, so $B$ is a group of exponent $2$ with neutral element 0.
@@
so $B$ is generated by the $q(a_{i},a_{j})$. By symmetry (Q2), we can always assume $i\leq j$ here. Furthermore, choosing $a_{0}=-1$ if $1\not=1\in G$, for any $i\in\{0,\ldots,n-1\}$, we get (cf. Remark 1.1)
@@
Hence, $B$ is generated by $(q(a_{i},a_{j})\mid 0\leq i<j\leq n-1)$ and, if $1\not=1$ additionally $q(a_{0},a_{0})$. The assertion follows. ∎
@@
Let S=(G,1,q)S=(G,-1,q) and $S^{\prime}=(G^{\prime},-1^{\prime},q^{\prime})$ be quaternionic structures with full 2B-groups $B$ and $B^{\prime}$. The following hold:
@@
If $F$ is a field and $I$ the fundamental ideal of its Witt ring, the factor group $I^{2}/I^{3}$ is isomorphic to the $2$-torsion part of the Brauer group by Merkurjev’s Theorem [Mer81]. In our abstract setting, the full $2B$-group will function as a substitute for $I^{2}/I^{3}$. We do so to avoid assuming $AP_{3}$ as an additional property of quaternionic structures. In the remainder of this section, we show that both $I^{2}/I^{3}$ and the full 2B-group are in fact the same for elementary types.
@@
For the reader’s convenience, we recall Marshall’s results from [Mar80, proof of 5.7, 5.17] on how the Arason-Pfister property $AP_{k}$ behaves for direct products and group extensions. Note that the statement for the direct product is only proven there for $k\in\{1,2\}$ but the proof works in the same way for larger $k$.
@@ 4. Quaternionic Matrices
In the following, we only consider structures $S$ of finite order $n\in\mathbb{N}$. In this section, we establish a unique matrix $M_{S}\in\operatorname{Mat}_{n}(\mathbb{N}_{0})$ by which we can identify $S$. This will be the groundwork for our computations later on. We usually write S=(G,1,q)S=(G,-1,q) with set of quaternions $Q$ and full $2B$-group $B$. We call its dimension $\dim(S):=\dim(B)$, which is finite by Lemma 3.1(c), the dimension of $S$.
@@
For $n\in\mathbb{N}$, we often consider matrices $M=(M_{i,j})_{0\leq i,j\leq n-1}\in\operatorname{Mat}_{n}(\mathbb{N}_{0})$ where we start counting the indices for rows and columns at 0. We bestow the set $\operatorname{Mat}_{n}(\mathbb{N}_{0})$ with the lexicographical order $_{\text{lex}}$, applied row by row, and extending the usual order on $\mathbb{N}_{0}$. We simply say that a matrix is ’smaller/greater/…’ than another if it is lexicographically smaller/greater/…
@@
defines a group of exponent 22 $(\mathbb{N}_{0},\ast_{\mathbb{N}_{0}})$ where any set $\{0,1,\ldots,2^{n}-1\}$ defines a subgroup with basis $(2^{k}\mid 0\leq k\leq n-1)$. This is easily verified using linear algebra. We often simply write $\ast$ instead of $\ast_{\mathbb{N}_{0}}$ and define the completion of $M$ as the matrix
@@
We also define $(B(M),\overline{\ast})$ as the free group of exponent 22 $\langle\overline{M}_{I,J}\mid I,J\subset\{0,\ldots,n-1\}\rangle$ with the (linkage) relations
@@
We say that a basis $\mathcal{A}=(a_{0},\ldots,a_{n-1})$ of $G$ is a basis of S if either 1=1-1=1 or $a_{0}=-1$. Given a basis $\mathcal{B}=(q_{0},\ldots,q_{m-1})$ of $B$, we define the group isomorphism
@@
If $\mathcal{B}$ is fixed, we often identify $(B,\ast)$ with $(\{0,\ldots,2^{m}-1\},\ast)$, writing $q_{\mathcal{B}}(a,b)$ or – in abuse of notation – simply $q(a,b)$ instead of $\phi_{\mathcal{B}}(q(a,b))$.
@@
Note that the matrices defined as above are in general not the incidence matrices as studied by Kula (cf. [Kul91]). The explicit matrices given here are in fact the lexicographically minimal ones for the respective structures which we leave for the reader to check. The minimality will be of importance for defining a unique quaternionic matrix of $S$.
@@
    • if $S\cong\mathbb{L}_{2k,0}$: $M_{\mathcal{A},\mathcal{B}}$ is the matrix $E_{2k}\in M_{2k}(\mathbb{N}_{0})$ with entry $1$ on the anti-diagonal and entry 0 else.
@@
(b): The properties (b)(M2) and (b)(M4) follow directly from (Q2) and (Q4). For (b)(M3), it suffices to see that q(a,a)=q(a,1)q(a,a)=q(a,-1) for all $a\in G$ by Remark 1.1 and this equals 0 if 1=11=-1 in $G$. Since we have chosen $a_{0}=-1$ in the other case, (b)(M3) follows.
@@
Let $M\in\operatorname{Mat}_{n}(\mathbb{N}_{0})$ be a prequaternionic matrix of type $\epsilon\in\{0,1\}$. Let $G$ be a group of exponent $2$ with neutral element $1$ and basis $\mathcal{A}:=(a_{0},\ldots,a_{n-1})$. If $\epsilon=0$, choose $1:=1\in G$, else $1:=a_{0}$. Define $Q$ as the set of entries of $\overline{M}$, naturally containing 0. Then
@@
defines a quaternionic structure S(M)=(G,1,q)S(M)=(G,-1,q) with $s(S)=1$ if $\epsilon=0$ and $s(S)>1$ else. For any basis $\mathcal{B}$ of the full $2B$-group of $S(M)$ and $I,J\subset\{0,\ldots,n-1\}$, it satisfies
@@
(a): Write $\mathcal{A}$ and $\mathcal{A}^{\prime}$ for the bases of $S$ and $S^{\prime}$ as in Proposition 4.4. It is clear that the basis transformation $\mathcal{A}\mapsto\mathcal{A}^{\prime}$ defines an isomorphism of quaternionic structures as it preserves the (1)(-1)-element ($M,M^{\prime}$ are of the same type) and the zero quaternions.
(b): As before, this follows from the fact that the canonical isomorphism of groups of exponent $2$ preserves 1-1 and the zero quaternions. ∎
@@
$(\mathbb{L}_{1}\times({\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\mathbb{L}_{1}}\times{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbb{L}_{1}}){\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\Delta}){\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}\Delta}=:S=:(G,-1,q)$
with square class group $G=\{1,a_{1}\}\times{\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\{1,b_{0}\}}\times{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\{1,a_{2}\}}\times{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\{1,a_{3}\}}\times{\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}\{1,a_{4}\}}$ according to construction and coloring. Then $a_{0}:=-1=a_{1}{\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}b_{0}}{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}a_{2}}$. We compute $M:=M_{\mathcal{A}}(S)$ for the basis $\mathcal{A}=(a_{0},\ldots,a_{4})$. We have $q(a_{0},a)\not=0$ for all $a\in G\setminus\{1\}$ ($S$ is pythagorean), so all entries of the first row of $M$ must be independent, meaning that by (c)(M0), we have (apply Proposition 5.1 from the following section or a straightforward computation) the first row (1 2 4 8 16)(1\ 2\ 4\ 8\ 16). Furthermore, by the properties of the direct product $\mathbb{L}_{1}\times({\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\mathbb{L}_{1}}\times{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbb{L}_{1}}){\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\Delta}$, we get $q(a_{1},{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}a_{2}})=0=q(a_{1},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}a_{3}})$. Applying (b)(M2) and (b)(M3) there are $4$ more values to compute:
@@
$\dim(S)=1+(({\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}1}+{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}1})+{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}2})+{\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}4}=9.$
@@
Let $\epsilon\in\{0,1\}$. For any $n\in\mathbb{N}$, we have a bijection between the set of normal quaternionic matrices of type $\epsilon$ in $\operatorname{Mat}_{n}(\mathbb{N}_{0})$ and the isomorphy types of quaternionic structures $S$ of order $n$ with 1=1-1=1 (if $\epsilon=0$) or $1\not=1$ (if $\epsilon=1$), mapping $M$ to $S(M)$.
@@
Let $M\in\operatorname{Mat}_{n}(\mathbb{N}_{0})$ be a quaternionic matrix of type $\epsilon\in\{0,1\}$. We say that $M^{\prime}\in\operatorname{Mat}_{n}(\mathbb{N}_{0})$ is a simple transformation of $M$ if there are $\in\{0,\ldots,n-1\}, iji\not=j$ such that
@@
The simple transformations display basis transformations of quaternionic structures in the following way: Take a quaternionic structure S=(G,1,q)S=(G,-1,q) with basis $\mathcal{A}=(a_{0},\ldots,a_{n-1})$ and basis $\mathcal{B}$ of the full $2B$-group. As $G$ is an $\mathbb{F}_{2}$ vector space, we cannot scale $\mathcal{A}$ nontrivially, so if $\mathcal{A}^{\prime}$ is another basis of $S$, we can obtain it from $\mathcal{A}$ by an iteration of swapping elements $a_{i}\leftrightarrow a_{j}$ or replacing $a_{i}$ by $a_{i}a_{j}$. The following lemma shows that the matrix transformations defined above translate the effects of these basis transformations to the quaternionic matrices, going from $M_{\mathcal{A},\mathcal{B}}$ to $M_{\mathcal{A}^{\prime},\mathcal{B}}$ by an iteration of simple transformations. For the canonical choice of $\mathcal{B}$ (cf. Lemma 4.2(c)) in this setting, we have to consider reduced quaternionic matrices:
@@
(c): First, we point out that applying simple transformations $(i\leftrightarrow j)$ or $(i\mapsto i\ast j)$ just means swapping the $i$-th and $j$-th basis element in the induced basis of $S(M)$ or replacing $a_{i}$ by $a_{i}a_{j}$. Both are basis transformations of $S(M)$ as by the choice of $i,j$, the element 1-1 is fixed. In particular, simple transformations and by iteration also general transformations of $M$ are again quaternionic matrices of the same type that are isomorphic to $M$. By (a), reduction of quaternionic matrices also does not change the isomorphy type, so the ’if’ part of the statement holds.